the areo of cross-section of a pipe is 6.5sq.cm and water is pumped out of it at the rate of 45km/h.find in litres,the volume of water which flows out of the pipe in one minute.

Answer by mananth(16949)

if you convert 45km/h to cm/minute (45*1000*100/60)= 75,000 cm/minute
we get 75,000 cm/minute
area = 6.5 cm^2
Volume = area * L
6.5*75,000 =487,500 cu.cm
=487.5 liters